# GUI-R1: A Generalist R1-Style Vision-Language Action Model For GUI Agents

### v3

@@ 3.1 Preliminaries
We define the goal of GUI agents in high-level instructional tasks as understanding and executing low-level instructions to complete the high-level task $Q$, based on the current interface image $I$, and the execution history $H$. Formally, given the input $Q$, $I$, and $H$, the model generates a set of candidate responses $O=\{o_{1},o_{2},\dots,o_{N}\}$, where each response contains attributes of the predicted low-level action $o^{\text{act}}$, input text $o^{\text{text}}$, and input point $o^{\text{point}}$. Each response is evaluated using a unified action space reward function to compute its reward $\{r_{1},r_{2},\dots,r_{N}\}$. GRPO [10] is applied to estimate advantages and update the policy model under KL divergence constraints. The relative advantage $A_{i}$ of the $i$-th response is calculated as follows:
@@ 3.2 Verifiable Rewards in Unified Action Space
Format reward. Following previous work [20; 10; 6], we introduce format rewards during training to evaluate whether the generated output adheres to the expected structural format, including both syntactic and semantic validity. Specifically, format rewards guide the model to generate reasoning processes and final answers in a structured format, which play a critical role in self-learning and iterative improvement during reinforcement fine-tuning. The format reward templates used in training and inference are as follows, where ‘$<$think$>$’ represents the reasoning process and ‘$<$answer$>$’ represents the final answer.
Unified Action Space Prompt for Task Training and Inference You are GUI-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [complete, close/delete, press_home, click, press_back, type, select, scroll, enter]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[complete, close/delete, press_home, click, press_back, type, select, scroll, enter], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
Unified Action Space Prompt for Grounding Training and Inference You are GUI-R1, a reasoning GUI Agent Assistant. In this UI screenshot $\boldsymbol{<image>}$, I want you to continue executing the command $\boldsymbol{task}$, with the action history being $\boldsymbol{history}$. Please provide the action to perform (enumerate from [click]), the point where the cursor is moved to (integer) if a click is performed, and any input text required to complete the action. Output the thinking process in $<$$\boldsymbol{think}$$>$ $<$$/\boldsymbol{think}$$>$ tags, and the final answer in $<$$\boldsymbol{answer}$$>$ $<$$/\boldsymbol{answer}$$>$ tags as follows: $<$$\boldsymbol{think}$>>><<<$/\boldsymbol{think}$$>$ $<$$\boldsymbol{answer}$$>$[‘action’: enum[click], ‘point’: [x, y], ‘input_text’: ‘no input text [default]’]$<$$/\boldsymbol{answer}$$>$.
@@
Response reward. The final response reward is composed of format rewards and accuracy rewards, defined as:$R_{o}=\alpha R_{\text{f}}+\beta R_{\text{acc}},$ where $R_{\text{f}}$ represents the format reward, $R_{\text{acc}}$ represents the accuracy reward, and $\alpha$ and $\beta$ are weighting parameters respectively.
@@ 4.1 Implementation Details
Training and inference details. For supervised fune-tuning (SFT), we use the QwenVL2.5-3B/7B [4] model as the base model for experiments and employ the LLaMA Factory [25] framework for one epoch of training to avoid overfitting. For RFT, we use the EasyR1 [26] framework for training over nine epochs. During inference, to ensure fairness, we apply a unified and simple prompt across all comparison methods, and conduct experiments under zero-shot prompt configurations. All experiments are conducted using 8$\times$×NVIDIA A100-80G GPUs.

### v4

@@ 3.1 Preliminaries
We define the goal of GUI agents in high-level instructional tasks as understanding and executing low-level instructions to complete the high-level task $Q$, based on the current interface image II, and the execution history $H$. Formally, given the input $Q$, II, and $H$, the model generates a set of candidate responses $O=\{o_{1},o_{2},\dots,o_{N}\}$, where each response contains attributes of the predicted low-level action $o^{\text{act}}$, input text $o^{\text{text}}$, and input point $o^{\text{point}}$. Each response is evaluated using a unified action space reward function to compute its reward $\{r_{1},r_{2},\dots,r_{N}\}$. GRPO [10] is applied to estimate advantages and update the policy model under KL divergence constraints. The relative advantage $A_{i}$ of the $i$-th response is calculated as follows:
@@ 3.2 Verifiable Rewards in Unified Action Space
Format reward. Following previous work [20; 10; 6], we introduce format rewards during training to evaluate whether the generated output adheres to the expected structural format, including both syntactic and semantic validity. Specifically, format rewards guide the model to generate reasoning processes and final answers in a structured format, which play a critical role in self-learning and iterative improvement during reinforcement fine-tuning. The format reward templates used in training and inference are as follows, where ‘<<think>>’ represents the reasoning process and ‘<<answer>>’ represents the final answer.
[…]
@@
[…]
@@ 4.1 Implementation Details
[…]